1 ELECTRIC CURRENT (10 PoINTS) An electrochemical cell consists of two silver electrodes placed in aqueous solution of silver nitrate. The positive electrode is called anode, whereas the negative electrode is called cathode. A constant current of 0.5 A is passed through the cell for 1 hour. a) What is the total charge g transported through the cell? b) If one electron is needed to discharge one positively charged silver ion (that is, to neutralize the ion and deposit it on the cathode), how many ions are discharged? c) What is the total mass of the deposited silver? (The mass of one silver atom is 107.9 u where u 1.66 x 10 27 kg is the atomic mass unit).
Answer #1


a) total charge is q = i*t = 0.5*1*60*60 = 1800 C


b) each electron has a charge 1.6*10^-19 C


then no.of ions required is 1800/(1.6*10^-19) = 1.125*10^22

C) mass of silver atom is 107.9*1.66*10^-27 kg

total mass is m = 1.125*10^22*107.9*1.66*10^-27


m = 2.01*10^-3 Kg